# Repeating nuclear transients from repeating partial tidal disruption events: reproducing ASASSN-14ko and AT2020vdq

### v1

@@ 3.1 Simulation Setup
3.2 Results for a $3M_{\odot}$ TAMS star
One of our primary goals in this work is to test the viability of the rpTDE model in providing an explanation for the repeated flares observed from ASASSN-14ko. To do this, we describe here the results of simulating the repeated partial disruption of a $3M_{\odot}$ TAMS star by a $10^{6}M_{\odot}$ SMBH, with an impact parameter $\beta=1$. The impact paramater $\beta\equiv r_{\rm t}/r_{\rm p}$ quantifies the strength of the encounter between the star and the SMBH. For a high-mass star at its late evolutionary stages, the analytical prediction for the critical impact parameter $\beta_{\rm c}$ required for its complete disruption exceeds this value by an appreciable amount (Coughlin & Nixon 2022; from Figure 4 of Bandopadhyay et al. (2024) a $3M_{\odot}$ TAMS star has a critical impact parameter $\beta_{\rm c}\sim 6$, though we note that general relativistic effects may reduce this value somewhat for very massive SMBHs; e.g., Beloborodov et al. 1992; Kesden 2012; Gafton et al. 2015; Tejeda et al. 2017; Gafton & Rosswog 2019; Ryu et al. 2020; Jankovič & Gomboc 2023). Thus, $\beta=1$ results in only a small amount of mass lost, and the star should – subject to the influence of the imparted rotation discussed above – be able to survive for many encounters while simultaneously fueling accretion events.
Each successive pericenter passage spins up the tidally stripped core by a small fraction of its break-up velocity. Figure 4 shows the velocity of the fluid comprising the core in its COM coordinates (i.e., both COM position and velocity removed), which qualitatively demonstrates that the core is indeed rotating following its tidal interaction with the SMBH. In the top panel of Figure 5, we show the angular velocity $\Omega$ imparted to the core during the first fifteen pericenter passages as a function of cylindrical radius $s$ within the core. The primary $y-$axis shows the angular velocity normalized by the break-up angular velocity of the core, $\Omega_{\rm c}=\sqrt{GM_{\rm c}/R_{\rm c}^{3}}$. Here, $R_{\rm c}$ is the core radius for a $3M_{\odot}$ TAMS star, defined as the radial distance from the center of the star at which the self-gravitational field of the star is maximized (Coughlin & Nixon, 2022), and $M_{\rm c}$ is the mass contained within the core. The secondary $y-$axis shows the angular velocity normalized by the angular velocity of the COM at pericenter, $\Omega_{\rm p}=\sqrt{2GM_{\bullet}/r_{\rm p}^{3}}$. This figure demonstrates that while the rotation rate is comparable to $\Omega_{\rm p}$, the imparted angular velocity is a small fraction of the break-up angular velocity of the core. The core is thus relatively unmodified, and this allows the bulk of the star to survive many encounters as it tidally interacts with the SMBH. The bottom panel of Figure 5 shows the corresponding linear velocity $v_{\varphi}\equiv s\Omega$, normalized by $\sqrt{GM_{\rm c}/R_{\rm c}}$.
@@ 3.2 Results for a $3M_{\odot}$ TAMS star
The amount of mass stripped during each encounter is $\mathcal{O}\left(0.01M_{\odot}\right)$, and gradually decreases over time, as shown in Figure 7. Roughly half of the tidally stripped stellar debris returns to the SMBH and accretes onto it. The fallback rates of stellar debris onto the SMBH for the first fifteen encounters are shown in Figure 8. The fallback rates are calculated using the binning procedure described in, e.g., Golightly et al. (2019b); Miles et al. (2020): at early times and when the particle flux is relatively large we calculate the fallback rate using equal time bins, whereas the late time fallback rate is binned by particle number. All of the fallback rates follow the late-time temporal scaling of $t^{-9/4}$, conforming to the expected scaling for partials (Coughlin & Nixon, 2019). Owing to the small amount of mass stripped in these encounters, the fallback rates obtained from the $10^{6}$ particle simulations exhibit noisy behavior. To vet the accuracy of the fallback rates obtained, we performed higher-resolution simulations using $10^{7}$ particles for the first four encounters. The fallback rates for the high-resolution simulations are compared against their low resolution counterparts in Figure 9. In general, we find good agreement between the fallback rates for corresponding encounters at different resolutions, with the most noticeable differences arising in the return time of the most-bound debris; because the outer layers of the star are better resolved at higher resolution, and the outer layers of the star coincide with the most-bound debris, the high-resolution simulations generally display earlier return times. The peak timescale decreases gradually from $\sim 80$ days to $\sim 45$ days over the first few encounters, and settles to an almost constant value of $\sim 45$ days beyond the first five encounters. The magnitude of the peak of the fallback rates is nearly constant, at $\sim 0.04M_{\odot}\rm yr^{-1}$, over all of the fifteen encounters shown in Figure 8.
@@ 3.3.2 $1M_{\odot}$ ZAMS star
Figure 13: Top: The amount of mass stripped $\Delta M$, as a function of the number of pericenter passages $N$, for a $1M_{\odot}$ ZAMS star with $\beta=0.6$ for each encounter. The absence of a centrally concentrated core-like structure leads to a progressive increase in the amount of mass stripped with the number of pericenter passages. Bottom: Fallback rates for the first five encounters between the $1M_{\odot}$ ZAMS star and a $10^{6}M_{\odot}$ SMBH, at $\beta=0.6$, with a resolution of $10^{6}$ particles. The peak of the fallback rate increases gradually as the amount of mass stripped increases, and the peak timescale shifts to earlier times.
The density profile of the $1M_{\odot}$ ZAMS star is not as centrally concentrated as that of a high-mass and evolved star. Consequently, the critical impact parameter $\beta_{\rm c}$ required for the complete disruption of the star is $\sim 1.8$ – which is approximately a factor of 3 smaller than that required to completely destroy a $3M_{\odot}$ TAMS star through tidal interactions with a $10^{6}M_{\odot}$ SMBH. Coughlin & Nixon (2022) estimated the minimum value of the impact parameter $\beta$ at which a star begins to lose any mass to be $\beta_{\rm partial}\simeq 0.6$, independent of the stellar properties. To simulate a grazing encounter between a $1M_{\odot}$ ZAMS star and a $10^{6}M_{\odot}$ SMBH, we chose an impact parameter $\beta=0.6$. The top panel of Figure 13 shows the amount of mass stripped in these simulations, as a function of the number of pericenter passages. For this star, the lack of a centrally concentrated core-envelope structure leads to an increasing amount of mass being stripped at each successive pericenter passage. The bottom panel of Figure 13 shows the fallback rates for the first five encounters. As seen in the figure, the peak timescale decreases as the number of pericenter passages increases. The magnitude of the peak of the fallback rate gradually increases, which can be attributed to the increase in the amount of mass lost as well as the decrease in the peak timescale. This result shows that young low-mass stars that are on grazing orbits around an SMBH are unlikely to give rise to repeated flares of comparable peak luminosity, as observed for ASASSN-14ko.
@@
Finally, to explore the other end of the spectrum of $\beta$ values pertinent to partials, we simulated two successive disruptions of the $1M_{\odot}$ ZAMS star at $\beta=1.5$, for which the fallback rates are shown in Figure 14. As noted above, the critical impact parameter for the complete disruption of this star is $\beta_{\rm c}\approx 1.8$ (Golightly et al., 2019b; Coughlin & Nixon, 2022), and thus for $\beta=1.5$, the star loses a significant fraction of its mass during the first pericenter passage. From the figure, we see that the late-time scaling of the fallback rate for the first encounter scales as $\propto t^{-9/4}$, indicating that the disruption is partial. However, the star is completely destroyed on its second pericenter passage, yielding a fallback rate that has a lower peak magnitude relative to the first, and that scales as $\propto t^{-5/3}$ at late-times. The dimmer second peak is consistent with the observed lightcurve for AT2018fyk (Wevers et al., 2023). Since the second encounter results in a complete disruption of the star, if AT2018fyk was generated by a low-mass star on an orbit having a $\beta$ value close to that required for its complete disruption, we would not expect to observe any more flares from this source.
@@ 4.1 Spin-up of the stellar core
In Section 2, we demonstrated using a simple toy model of a dumbbell (composed of two point masses connected by a rigid rod) being gravitationally torqued through tidal interactions with a black hole, that once the dumbbell is spun up to a rotational velocity $\Omega_{\rm p}=\sqrt{GM_{\bullet}}/r_{\rm p}^{3/2}$, the angular velocity imparted to the dumbbell during subsequent encounters with the black hole does not increase monotonically, and is instead an oscillatory function of time about a mean value of $\sim\Omega_{\rm p}$. While this toy model demonstrates that tidal interactions with an SMBH would not result in a star being spun up to its break-up rotational velocity, it does not capture all aspects of the gas-dynamic response of a gravitationally bound star that interacts tidally with an SMBH. In particular, and as seen in the results of the hydrodynamical simulations presented in Section 3, the tidal interaction leads to a secular increase in the angular velocity imparted to the core, as the increase in the size of the core implies that the star is more efficiently gravitationally-torqued by the SMBH, and a renewed quadrupole moment is excited on each encounter, the nonlinear interaction of which with the tidal field of the SMBH is ultimately responsible for generating the net rotation. However, our results for the $3M_{\odot}$ TAMS star on a $\beta=1$ orbit show that the angular velocity imparted to the surviving stellar core over multiple encounters with an SMBH is a small fraction of its break-up spin, and asymptotes to a value of $0.07\Omega_{\rm c}$ in the limit of a large number of pericenter passages.
@@
Finally, the rotation imparted to the star is a nonlinear effect: the tidal bulge provides the quadrupole moment, which subsequently experiences the torque applied by the tidal field. Consequently, we would expect this effect to decline extremely strongly with distance from the SMBH, similar to the way in which the energy imparted via tides falls off with $\beta$ very rapidly (Press & Teukolsky, 1977). Thus, the statement that the star is spun up to $\sim\Omega_{\rm p}$ is likely only valid for $0.5\lesssim\beta\lesssim\beta_{\rm c}$, where $\beta_{\rm c}$ is the value of $\beta$ at which the star is completely destroyed.
@@ 4.2 Effect of Stellar Structure on Survivability
Our hydrodynamical simulations show that the survivability of a star undergoing repeated tidal interactions with an SMBH depends on its structural properties. High-mass stars that are in their late evolutionary stages develop a core-envelope structure that enhances their probability of survival relative to low-mass, less centrally concentrated stars. This is in agreement with the results of  Liu et al. (2023), where the authors concluded, based on an adiabatic mass-loss model, that a centrally concentrated high-mass star having a diffuse outer envelope can generate ASASSN-14ko like flares as it undergoes periodic mass-loss events on a grazing orbit around an SMBH. We showed that the $3M_{\odot}$ TAMS star and the $1.3M_{\odot}$ TAMS star (for which the critical impact parameter required for their complete disruption is $\beta_{\rm c}\simeq 6$) on a $\beta=1$ orbit can survive many encounters, losing a small amount of mass on each pericenter passage, and generating luminous flares for a long time. On the other hand, for the $1M_{\odot}$ ZAMS star, the density profile is not as centrally concentrated, and the amount of mass stripped per encounter is a monotonically increasing function of the number of pericenter passages, which can be detrimental to the survival of the star.
@@
Antonini et al. (2011) used $N-$body integrations and SPH simulations to study the repeated tidal stripping of main-sequence of stars as one of the possible outcomes of the tidal break-up of binary star systems by SMBHs at the centers of galaxies. They used $N-$body integrations to evolve the binaries for $\gtrsim 100$ orbits around the SMBH, and found that in a Hills-capture scenario, where one of the stars is ejected as a hyper-velocity star, the bound companion can undergo periodic mass-loss episodes for hundreds of orbits before being completely destroyed. Their SPH simulations, which have a maximum resolution of $\sim 4\times 10^{4}$ particles, do not explore the parameter space of stellar and orbital parameters that we considered here, but our results qualitatively agree with their conclusion – that main-sequence stars on highly eccentric bound orbits around SMBHs can undergo repeated tidal disruptions, giving rise to periodic flares of varying intensity. We demonstrated that the strength, duration and periodicity of these flares are largely determined by the structural properties of the star and the distance of closest approach between the star and the SMBH.
@@
Zalamea et al. (2010) studied the feasibility of detection of electromagnetic and gravitational-wave signals arising from an extended period of slow mass-loss in white dwarfs on inspiralling orbits around a $10^{5}M_{\odot}$ SMBH. While the region of parameter space explored in their work does not directly correlate to ours, they found that white dwarfs on highly eccentric orbits around an SMBH, undergoing slow periodic mass-loss episodes can last for thousands of orbits, emitting gravitational-waves detectable in the Laser Interferometer Space Antenna (LISA) band, and accretion flares with luminosities close to the Eddington limit of the SMBH. MacLeod et al. (2013) explored a different region of the parameter space for repeated mass-transfer and accretion onto SMBHs through tidal interactions with bound stars. They used a combination of hydrodynamical simulations and a semi-analytic model to study the tidal stripping of a giant star with mass $M_{\star}=1.4M_{\odot}$ and $R_{\star}=50R_{\odot}$, on an eccentric orbit around a $10^{7}M_{\odot}$ SMBH. The mass-transfer process is initiated when the star evolves off its main-sequence phase, and up the red giant branch, as its mean density drops, making it susceptible to mass-loss at pericenter. They showed that the red-giant star can survive for hundreds of orbits around the SMBH, giving rise to low-intensity flares that recur on the orbital timescale of the star, which is $\gtrsim 10^{3}$ years.
@@ 4.3 Evolution of the Orbital Period and the Peak Fallback Timescale
The secular evolution of the orbital period of ASASSN-14ko is not well understood, and various mechanisms have been suggested for reproducing the observed period derivative of $\dot{P}=-0.0016$. Payne et al. (2021) and Cufari et al. (2022a) concluded that the change in orbital period due to the emission of gravitational waves underestimates the period derivative by several orders of magnitude. Linial & Quataert (2024) analyzed various physical processes that could be responsible for producing the observed period change in ASASSN-14ko. They suggest that the hydrodynamical drag experienced by the star as it interacts with the accretion disc of the SMBH provides the most likely explanation for the observed change in period. However, they note that the minimum disc mass required for this likely exceeds the mass of the star, thereby necessitating the existence of an AGN-like disc to explain the observed evolution of the period.
@@
Our analytical estimates from Section 2 show that for solar-like stars being repeatedly tidally disrupted by an SMBH, the angular velocity imparted to the stellar core can change the orbital period by a small fraction of its original Keplerian time period. For the specific case of ASASSN-14ko, having a period of $114\pm 1$ days, we estimated the fractional change in the time period to be $\Delta T/T\simeq-0.001$, which is in excellent agreement with the observed period derivative for this event. We note, however, that to continue to reduce the orbit by the same fractional amount, this energy must be lost from the star efficiently. Liu et al. (2024) argued that the rapid evolution of the recurrence time of the flares observed from the eRASSt-J045650 source, which has now flared $\sim 5$ times, can be interpreted as a change in the orbital period of a $1M_{\odot}$ star undergoing repeated partial disruptions on an orbit around a $10^{5}M_{\odot}$ SMBH. However, they estimated that the amount of mass that the star must lose on a single pericenter passage is $\sim 0.8-0.9M_{\odot}$. Given that this is a substantial fraction of the original mass of the star, it seems unlikely that the remnant of this interaction would survive subsequent pericenter passages without undergoing a drastic and unphysical change in its pericenter distance. Additionally, as argued in Cufari et al. (2022a) and as shown here in Section 2, reducing the period by the amount claimed in Liu et al. (2024) would necessitate imparting more energy into the star than its own binding energy. While the change in orbital period induced by the spin-up of the stellar core can only be an extremely small fraction of the original period, the peak timescale for the fallback rate of tidally stripped stellar debris can exhibit greater variation due to the imparted rotation. We found from the simulation of the partial disruption of a $3M_{\odot}$ TAMS star by a $10^{6}M_{\odot}$ SMBH, with an impact parameter $\beta=1$, that the peak fallback time scales with the imparted angular velocity $\Omega$ approximately as $\left(1+\sqrt{2}\Omega/\sqrt{GM_{\bullet}/r_{\rm p}^{3}}\right)^{-0.8}$. From Figure 10 we see that if an rpTDE is observed during the first few pericenter passages of the star, the peak timescale could show significant variation between successive flares, but as the angular momentum imparted to the stellar core converges to a nearly constant value over multiple encounters, the variation in the peak timescale would be less prominent for later pericenter passages.

### v2

@@ 3.2 Results for a $3M_{\odot}$ TAMS star
The amount of mass stripped during each encounter is $\mathcal{O}\left(0.01M_{\odot}\right)$, and gradually decreases over time, as shown in Figure 7. The fallback rates of stellar debris [^4] In our simulations the star is approximated to be on a parabolic orbit, in which case only half of the mass that is tidally stripped from the star returns to the SMBH, with the other half unbound. Because the star must actually be on an elliptical orbit, a fraction of the other (usually unbound) tail can be bound to the SMBH, and will thereby produce a second, relatively low-level (but potentially observable) accretion event; the amount of mass in this “less bound” tail depends on the eccentricity of the orbit, the impact parameter $\beta$, and the SMBH mass (Hayasaki et al., 2018; Park & Hayasaki, 2020; Cufari et al., 2022b). onto the SMBH for the first fifteen encounters are shown in Figure 8. The fallback rates are calculated using the binning procedure described in, e.g., Golightly et al. (2019b); Miles et al. (2020): at early times and when the particle flux is relatively large we calculate the fallback rate using equal time bins, whereas the late time fallback rate is binned by particle number. All of the fallback rates follow the late-time temporal scaling of $t^{-9/4}$, conforming to the expected scaling for partials (Coughlin & Nixon, 2019). Owing to the small amount of mass stripped in these encounters, the fallback rates obtained from the $10^{6}$ particle simulations exhibit noisy behavior. To vet the accuracy of the fallback rates obtained, we performed higher-resolution simulations using $10^{7}$ particles for the first four encounters. The fallback rates for the high-resolution simulations are compared against their low resolution counterparts in Figure 9. In general, we find good agreement between the fallback rates for corresponding encounters at different resolutions, with the most noticeable differences arising in the return time of the most-bound debris; because the outer layers of the star are better resolved at higher resolution, and the outer layers of the star coincide with the most-bound debris, the high-resolution simulations generally display earlier return times. The peak timescale decreases gradually from $\sim 80$ days to $\sim 45$ days over the first few encounters, and settles to an almost constant value of $\sim 45$ days beyond the first five encounters. The magnitude of the peak of the fallback rates is nearly constant, at $\sim 0.04M_{\odot}\rm yr^{-1}$, over all of the fifteen encounters shown in Figure 8.
@@ 3.3.2 $1M_{\odot}$ ZAMS star
Figure 13: Top: The amount of mass stripped, $\Delta M$, as a function of the number of pericenter passages, $N$, for a $1M_{\odot}$ ZAMS star with $\beta=0.6$ for each encounter. The absence of a centrally concentrated core-like structure leads to a progressive increase in the amount of mass stripped with the number of pericenter passages. Bottom: Fallback rates for the first five encounters between the $1M_{\odot}$ ZAMS star and a $10^{6}M_{\odot}$ SMBH, at $\beta=0.6$, with a resolution of $10^{6}$ particles. The peak of the fallback rate increases gradually as the amount of mass stripped increases, and the peak timescale shifts to earlier times.
The density profile of the $1M_{\odot}$ ZAMS star is not as centrally concentrated as that of a high-mass and evolved star. Consequently, the critical impact parameter $\beta_{\rm c}$ required for the complete disruption of the star is $\sim 1.8$ – which is approximately a factor of 3 smaller than that required to completely destroy a $3M_{\odot}$ TAMS star through tidal interactions with a $10^{6}M_{\odot}$ SMBH. Coughlin & Nixon (2022a) estimated the minimum value of the impact parameter $\beta$ at which a star begins to lose any mass to be $\beta_{\rm partial}\simeq 0.6$, independent of the stellar properties. To simulate a grazing encounter between a $1M_{\odot}$ ZAMS star and a $10^{6}M_{\odot}$ SMBH, we chose an impact parameter $\beta=0.6$. The top panel of Figure 13 shows the amount of mass stripped in these simulations as a function of the number of pericenter passages. For this star, the lack of a centrally concentrated core-envelope structure leads to an increasing amount of mass being stripped at each successive pericenter passage. The bottom panel of Figure 13 shows the fallback rates for the first five encounters. As seen in the figure, the peak timescale decreases as the number of pericenter passages increases. The magnitude of the peak of the fallback rate gradually increases, which can be attributed to the increase in the amount of mass lost as well as the decrease in the peak timescale. This result shows that young low-mass stars that are on grazing orbits around an SMBH are unlikely to give rise to repeated flares of comparable peak luminosity, as observed for ASASSN-14ko.
@@
Finally, to explore the other end of the spectrum of $\beta$ values pertinent to partials, we simulated two successive disruptions of the $1M_{\odot}$ ZAMS star at $\beta=1.5$, for which the fallback rates are shown in Figure 14. As noted above, the critical impact parameter for the complete disruption of this star is $\beta_{\rm c}\approx 1.8$ (Golightly et al., 2019b; Coughlin & Nixon, 2022a), and thus for $\beta=1.5$, the star loses a significant fraction of its mass during the first pericenter passage. From the figure, we see that the late-time scaling of the fallback rate for the first encounter scales as $\propto t^{-9/4}$, indicating that the disruption is partial. However, the star is completely destroyed on its second pericenter passage, yielding a fallback rate that has a lower peak magnitude relative to the first, and that scales as $\propto t^{-5/3}$ at late-times. The dimmer second peak is consistent with the observed lightcurve for AT2018fyk (Wevers et al., 2023). Since the second encounter results in a complete disruption of the star, if AT2018fyk was generated by a low-mass star on an orbit having a $\beta$ value close to that required for its complete disruption, we would not expect to observe any more flares from this source. However, the detection of a second prompt shutoff implies that this explanation likely fails for AT2018fyk, and suggests a third rebrightening in 2027 (Pasham et al., 2024a).
@@ 4.1 Spin-up of the stellar core
In Section 2 we used a simple toy model of a “dumbbell” (composed of two point masses connected by a rigid rod) gravitationally interacting with a black hole to argue that a star should not be torqued beyond a rotational velocity of $\Omega_{\rm p}=\sqrt{GM_{\bullet}}/r_{\rm p}^{3/2}$. While this outcome and toy model agreed qualitatively with our hydrodynamical simulations, there were important differences. Specifically, instead of being torqued to $\sim\Omega_{\rm p}$ on the first encounter with a pseudo-random oscillation in its rotational velocity thereafter, the star was imparted a small change (which declined with the number of pericenter passages; see Figure 5) in $\Omega$ each time it interacted strongly and tidally with the SMBH. This can be attributed to both an increase in the size of the star, such that the star is more efficiently gravitationally torqued by the SMBH, and the fact that a renewed quadrupole moment is excited on each encounter, the nonlinear interaction of which with the tidal field of the SMBH is ultimately and at least partially responsible for generating the net rotation. Our results for the $3M_{\odot}$ TAMS star on a $\beta=1$ orbit show that the angular velocity imparted to the surviving stellar core over multiple encounters with an SMBH is a small fraction of its break-up spin, and asymptotes to a value of $0.07\Omega_{\rm c}$ in the limit of a large number of pericenter passages.
@@
Finally, the rotation imparted to the star is a nonlinear effect: the tidal bulge provides the quadrupole moment, which subsequently experiences the torque applied by the tidal field. Additionally, Kochanek (1992) argued that the tidal interaction excites modes that have non-zero vorticity, and the damping of these modes establishes a bulk rotation within the star, such that the angular momentum in the oscillatory modes and that in the rotation of the star cancel out the net vorticity. We discuss the development of solid-body rotation within the star in Appendix B and show that it is insensitive to numerical resolution, i.e., it is a physical consequence of the tidal interaction with the SMBH. Because it is a nonlinear effect, we would expect the net rotation per encounter to decline extremely strongly with distance from the SMBH, similar to the way in which the energy imparted via tides falls off with $\beta$ very rapidly (Press & Teukolsky, 1977). Thus, the statement that the star is spun up to $\sim\Omega_{\rm p}$ is likely only valid for $0.5\lesssim\beta\lesssim\beta_{\rm c}$, where $\beta_{\rm c}$ is the value of $\beta$ at which the star is completely destroyed.
@@ 4.2 Effect of Stellar Structure on Survivability
[…]
@@
Antonini et al. (2011) used $N-$body integrations and SPH simulations to study the repeated tidal stripping of main-sequence of stars as one of the possible outcomes of the tidal break-up of binary star systems by SMBHs. They used $N-$body integrations to evolve the binaries for $\gtrsim 100$ orbits around the SMBH, and found that in a Hills-capture scenario, where one of the stars is ejected as a hyper-velocity star, the bound companion can undergo periodic mass-loss episodes for hundreds of orbits before being completely destroyed. Their SPH simulations, which have a maximum resolution of $\sim 4\times 10^{4}$ particles, do not explore the parameter space of stellar and orbital parameters that we considered here, but our results qualitatively agree with their conclusion – that main-sequence stars on highly eccentric bound orbits around SMBHs can undergo repeated tidal disruptions, giving rise to periodic flares of varying intensity. We demonstrated that the strength, duration and periodicity of these flares are largely determined by the structural properties of the star and the distance of closest approach between the star and the SMBH.
@@
Zalamea et al. (2010) studied the feasibility of detecting electromagnetic and gravitational-wave signals arising from an extended period of slow mass-loss in white dwarfs on inspiralling orbits around a $10^{5}M_{\odot}$ SMBH. While the region of parameter space explored in their work does not directly correlate to ours, they found that white dwarfs on highly eccentric orbits around an SMBH, undergoing slow periodic mass-loss episodes, can last for thousands of orbits, emitting gravitational-waves detectable in the Laser Interferometer Space Antenna (LISA) band and accretion flares with luminosities close to the Eddington limit of the SMBH. MacLeod et al. (2013) used a combination of hydrodynamical simulations and a semi-analytic model to study the tidal stripping of a giant star with mass $M_{\star}=1.4M_{\odot}$ and $R_{\star}=50R_{\odot}$, on an eccentric orbit around a $10^{7}M_{\odot}$ SMBH. The mass-transfer process is initiated when the star evolves up the red giant branch. They showed that the red-giant star can survive for hundreds of orbits around the SMBH, giving rise to low-intensity flares that recur on the orbital timescale of the star, which is $\gtrsim 10^{3}$ years.
@@ 4.3 Evolution of the Orbital Period and the Peak Fallback Timescale
The secular evolution of the orbital period of ASASSN-14ko is not well understood, and various mechanisms have been suggested for reproducing the observed period derivative of $\dot{P}=-0.0026$. Payne et al. (2021) and Cufari et al. (2022a) concluded that gravitational wave emission underestimates the period derivative by several orders of magnitude. Linial & Quataert (2024) analyzed various physical processes that could be responsible for producing the observed period change in ASASSN-14ko, and suggested that the hydrodynamical drag experienced by the star as it interacts with the accretion disc of the SMBH provides the most likely explanation for the observed change in period. However, they noted that the minimum disc mass required for this likely exceeds the mass of the star, thereby necessitating the existence of an AGN-like disc to explain the observed evolution of the period.
Our analytical estimates from Section 2 show that for solar-like stars being repeatedly tidally disrupted by an SMBH, the angular velocity imparted to the stellar core can change the period by a small fraction of its original value. For the specific case of ASASSN-14ko, having a period of $114\pm 1$ days, we estimated the fractional change in the period to be $\Delta T/T\simeq-0.003$, which is in excellent agreement with observations (Payne et al., 2023). We note, however, that to continue to reduce the orbit by the same fractional amount, this energy must be lost from the star and efficiently, i.e., on a per-orbit basis. Since the imparted angular momentum to the star results in an velocity that is approximately uniform, it is unlikely that this energy will be dissipated efficiently, meaning that – if the imparted rotation does represent the sink of orbital energy that leads to the period decay – one would expect a pronounced $\ddot{P}$ as the star is not efficiently torqued above $\Omega_{\rm p}$.
In addition to bulk rotation, one expects a comparable amount of energy produced in oscillatory modes that is ultimately dissipated as heat, either viscously or through three-mode (and higher) nonlinear couplings (e.g., McMillan et al. 1987; Kochanek 1992; Kumar & Goodman 1996; Weinberg et al. 2012). As we discuss in Appendix A, we expect most of the thermal energy to be concentrated in the outer layers of the star where the amplitude of the tidal acceleration is largest, and these layers are removed on subsequent encounters with the SMBH. For ASASSN-14ko (i.e., with an orbital period of 114 days and a SMBH mass of $10^{7}M_{\odot}$), the reduction in the orbital period would result in the deposition of $\sim 1.5\times 10^{46}$ erg in $0.01M_{\odot}$ (i.e., comparable to the amount of mass stripped from the star in our simulations), which would require a luminosity of $L\approx 3.5L_{\rm Edd}$ if this energy were to be lost radiatively ($L_{\rm Edd}$ is the Eddington luminosity of a $3M_{\odot}$ star, assuming $\kappa_{\rm es}=0.34$ cm2 g-1 for the opacity; see the discussion in Appendix A for more details). The amount of energy dissipated via tides and required to reproduce the observed $\dot{P}$ therefore could not be efficiently exhausted radiatively (see also Kumar & Quataert 1996), but the tidal stripping of the outer layers of the star serves an alternative outlet through which the energy can be lost mechanically and without destroying the high-density (and thermally insulated; again, see Appendix A) core, thus allowing periodic accretion events to continue. While our resolution tests and the implementation of alternative thermodynamic treatments, discussed in detail in Appendix A, provide evidence to suggest that this is the means by which the star can continue to survive many tidal encounters with the SMBH, additional investigations – particularly in the context of linear tidal theory – are required to solidify this connection.
@@
Liu et al. (2024b) argued that the rapid evolution of the recurrence time of the flares observed from eRASSt-J045650, which has now flared $\sim 5$ times, can be interpreted as a change in the orbital period of a $1M_{\odot}$ star undergoing repeated partial disruptions on an orbit around a $10^{5}M_{\odot}$ SMBH. However, they estimated that the amount of mass that the star must lose on a single pericenter passage is $\sim 0.8-0.9M_{\odot}$. Given that this is a substantial fraction of the original mass of the star, it seems unlikely that the remnant of this interaction would survive subsequent pericenter passages without undergoing a drastic and unphysical change in its pericenter distance. Additionally, as argued in Cufari et al. (2022a) and as shown here in Section 2, reducing the period by the amount claimed in Liu et al. (2024b) would necessitate imparting more energy into the star than its own binding energy.
While the change in orbital period induced by the spin-up of the stellar core can only be an extremely small fraction of the original period, the peak timescale for the fallback rate of tidally stripped stellar debris can exhibit greater variation due to the imparted rotation. We found from the simulation of the partial disruption of a $3M_{\odot}$ TAMS star by a $10^{6}M_{\odot}$ SMBH that the peak fallback time scales with the imparted angular velocity $\Omega$ approximately as $\left(1+\sqrt{2}\Omega/\sqrt{GM_{\bullet}/r_{\rm p}^{3}}\right)^{-0.8}$. From Figure 10 we see that if an rpTDE is observed during the first few pericenter passages of the star, the peak timescale could show significant variation between successive flares, but as the angular momentum imparted to the stellar core converges to a nearly constant value over multiple encounters, the variation in the peak timescale would be less prominent for later pericenter passages.